### v1

# Near-horizon polarized images of rotating hairy Horndeski black holes

Abstract
Recently, Hou et al. [Astrophys. J. Lett. 988, L51 (2025)] revealed that the EVPA (Electric Vector Position Angle) of polarization vectors in the near-horizon images is governed solely by spacetime geometry and is irrespective of the plasma flows. Here, we generalize the study to the scenario of a rotating hair black hole within Horndeski gravity and probe the effects of the hairy parameter on the EVPA. For a fixed inclination, the hairy parameter enhances the observed EVPA in the slowly rotating case, but decreases it in the rapidly rotating case. For a fixed black hole spin, the influence of the hairy parameter on the observed EVPA under different observer inclinations is further modulated by the azimuthal angle of the observed polarization vector. The hairy parameter’s impact is more distinct in the low inclination case as the azimuthal angle lies within a specific range, but is almost independent of the observer inclination as the azimuthal angle is beyond this specific range. Furthermore, the dependence of the hairy parameter’s impact on the EVPA is stronger with respect to the black hole spin than to the inclination angle. These results could help to further understand the near-horizon polarized images and Horndeski gravity.
pacs:
04.70.Bw, 04.25.-g, 97.60.Lf
I Introduction
The images of supermassive black holes released by the Event Horizon Telescope (EHT) collaboration EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019pgp ; EventHorizonTelescope:2019ggy ; EventHorizonTelescope:2021srq ; EventHorizonTelescope:2022wkp ; EventHorizonTelescope:2022urf provide compelling evidence for the existence of black holes. The polarization patterns stored in these images EventHorizonTelescope:2021bee ; EHT:2024nwx ; EHT:2024ehx reveal that there exists the magnetic field around supermassive black holes. Therefore, the polarized images of black holes are expected as novel avenues to explore the matter distribution surrounding black holes, to identify black hole parameters, as well as to test the theory of general relativity. It is well known that the observed polarization images heavily depend on the plasma accretion flows around black holes. Therefore, in order to accurately extract the information on black holes, it is crucial to disentangle the corresponding plasma effects from the images. Recently, the near-horizon polarization pattern is found to be determined solely by the spacetime geometry and observer’s inclination, while is irrespective of the plasma flows Hou:2024qqo . This means that the near-horizon polarization pattern provides a clean and precise probe of extracting the black hole information including the black hole spin and other fundamental parameters. Furthermore, their analysis indicates that such polarization pattern may be detectable with future high-resolution observations, such as EHT arrays and spacebased VLBI. However, their research focuses on only the Kerr and Kerr-Newman-Nut black hole spacetime. Thus, it is necessary to study near-horizon polarization patterns for other black holes within theories of the alternative gravity.
The simplest and non-trivial modifications of general relativity are probably scalar-tensor theories, which contain a scalar degree of freedom besides the gravitational metric. Horndeski gravity is one of the most famous theories scalar-tensor theories, which plays a significant role in explaining the accelerating expansion of current universe Kase:2018aps . Although Horndeski theory contains higher derivatives of a scalar field, it owns the diffeomorphism invariance and the second-order differential field equations. Another good behavior of Horndeski gravity is that there is no Ostrogradsky instabilitygw ; Deffayet:2011gz . Therefore, Horndeski theory provides a natural platform for verifying the no-hair theorem of black holes. The hairy black holes in Horndeski gravity have been widely studied including the radially dependent scalar field Rinaldi:2012vy ; Cisterna:2014nua ; Feng:2015wvb ; Sotiriou:2013qea ; Miao:2016aol ; Kuang:2016edj ; Babichev:2016rlq ; Benkel:2016rlz ; Filios:2018xvy ; Cisterna:2018hzf ; Giusti:2021sku ; Cisterna:2015uya ; Anabalon:2013oea ; Fontana:2018fof and the time-dependent scalar field Babichev:2013cya ; Babichev:2017lmw ; BenAchour:2018dap ; Takahashi:2019oxz ; Minamitsuji:2019shy ; Arkani-Hamed:2003juy . Recently, a static and spherical symmetric hairy black hole solution of Horndeski gravity has been obtained with the quartic scalar field model Bergliaffa:2021diw . This Horndeski hairy carries a hairy parameter with the dimension of length. The presence of the hairy parameter yields that the metric contains an additional log-term deforming from the standard Schwarzschild one Bergliaffa:2021diw . For this static hairy Horndeski black hole, studies of the strong gravitational lensing RKWalia ; Kumar:2021cyl ; Atamurotov:2022slw and the black hole image Shi:2024bpm ; Wang:2023vcv ; Ghosh:2022kit ; Vagnozzi:2022moj ; Afrin:2021wlj show that the Horndeski hair has a significant influence on the related observable. Moreover, a rotating black hole with Horndeski hair has been obtained recently by making use of the revised Newman–Janis algorithm RKWalia . The imprints of Horndeski hair together with black hole spin have been extensively studied in various theoretical and observational phenomena. For example, the hairy parameter makes the shadow of a black hole have a larger size and a greater degree of deformation. The presence of hair in the black hole enhances the superradiance and its frequency rangeJha:2022tdl , so the energy extraction within Horndeski framework is more efficient than in the pure Kerr case. Furthermore, effects of the hairy parameter on black hole stability under perturbations Lei:2023wlt ; Yang:2023lcm , accretion process Donmez:2024lfi , and precessing and periodic orbits Lin:2023eyd ; Lin:2023rmo have been studied in this rotating hairy black hole spacetime. These novel properties induced by the hairy parameter are helpful for understanding the gravitational structure of the specific theory.
In this paper, we will study the near-horizon polarized images of the rotating hairy black hole in Horndeski theory. Our study motivates from two aspects. One is to check whether the observed EVPA of electromagnetic emission from the near-horizon region is independent of the plasma flows and is determined only by spacetime geometry in the rotating hairy Horndeski black hole spacetime. The other is to probe how hairy parameter affect the observed EVPA in the near-horizon images, which is crucial to the issue whether the near-horizon polarized images provide a powerful way of detecting black hole hair, as well as examining theories of gravity.
This paper is organized as follows. In Sec. II, we present the observed EVPA for electromagnetic emission from the near-horizon region of the rotating hairy Horndeski black hole through analyzing the equatorial plasma flows with ideal conductivity as in Hou:2024qqo . In Sec.III, we analyze effects of the hairy parameter on the observed EVPA in the near-horizon images, and further compare with effects from the black hole spin and the observation inclination. Finally, we end the work with a summary.
II Observed polarization near the horizon of a rotating hairy Horndeski black hole
In this section, we will investigate the observed polarization in the near-horizon images of a rotating hair black hole within Horndeski gravity. Let us now to review briefly the rotating hairy Horndeski black hole, which is a stationary and spherically symmetric black hole solution with scalar hair. With the standard Boyer-Lindquist coordinates, its metric has a form RKWalia
$ds^{2}$ $=$ $-\left(\frac{\Delta-a^{2}\sin^{2}\theta}{\Sigma}\right)dt^{2}+\frac{\Sigma}{\Delta}{dr^{2}}+\Sigma d\theta^{2}+\frac{2a\sin^{2}\theta}{\Sigma}\left[\Delta-(r^{2}+a^{2})\right]dtd\varphi$ (1)
$+$ $\frac{\sin^{2}\theta}{\Sigma}\left[(r^{2}+a^{2})^{2}-\Delta a^{2}\sin^{2}\theta\right]d\varphi^{2},$
with
$\Delta=r^{2}+a^{2}-2Mr+hr\ln\left(\frac{r}{2M}\right),\quad\quad\quad\Sigma=r^{2}+a^{2}\cos^{2}\theta.$ (2)
Here $M$ and $a$ respectively correspond to the mass and spin parameters of the black hole, and the parameter $h$ is hair parameter with length dimension. The metric (1) reduces to the Kerr case as the parameter $h\rightarrow 0$, and to the static hairy Horndeski black hole Bergliaffa:2021diw as $a=0$. From the metric (1), one can find that the black hole horizons are related to the parameter $h$. Figure 1 shows the relation between the black hole horizon and the hair parameter $h$ for different black hole spins. We can find that for $h_{c}<h\leq 0$, the black hole has two horizons: the outer horizon $r_{+}$ and the inner horizon $r_{-}$. When $h=h_{c}$, the two horizons coincide and the black hole becomes an extremal one. Moreover, the critical value $h_{c}$ increases monotonically with the spin of the black hole, which is also shown in Fig.2.
Figure 1: Variation of the function $\Delta{(r)}$ with the radial coordinate $r$ for different black hole hair parameters. The black hole spins in the left and right panels are set to $a=0.2$ and $a=0.94$, respectively. Here set $M=1$.
Figure 2: Variation of the critical value $h_{c}$ with the spin $a$ for rotating hair black holes within Horndeski gravity. Here set $M=1$.
We are now in position to study the polarization vector of plasma near the horizon of hairy Horndeski black holes. As a main component in accretion disk around black holes, plasma is a state of matter composed of charged particles, including electrons and ions. Generally, the dynamics of plasma near black holes can be described by the general relativistic magnetohydrodynamics (GRMHD). In ref.Hou:2023bep , a simplified magnetohydrodynamic model has been constructed within GRMHD framework and the corresponding magnetic field of flows is expressed as Ruffini:1975ne ; Hou:2023bep
$B^{\mu}=\frac{\Psi}{\sqrt{-g}U^{r}}(U_{t}U^{\mu}+\delta_{t}^{\mu}),$ (3)
where $g$ is the determinant of the spacetime metric and $U^{\mu}$ is the plasma bulk velocity. The quantity $\Psi$ denotes the overall strength and is conserved along the bulk velocity. Since the millimeter-wavelength emission is dominated by the collective synchrotron radiation of thermal or non-thermal electrons, its polarization vector $f^{\mu}$ is largely perpendicular to the global magnetic field 1979Lightman . Therefore, the polarization vector $f^{\mu}$ of the plasma can be written as in a covariant form Hou:2024qqo
$f^{\mu}=\frac{\epsilon^{\mu\nu\alpha\beta}U_{\nu}p_{\alpha}B_{\beta}}{\omega\sqrt{B_{\bot}^{2}}}\,$, (4)
where $\varepsilon_{\mu\nu\alpha\beta}$ denotes the Levi-Civita tensor and $p^{\mu}$ represents the photon four-momentum. $\omega=-U^{\mu}p_{\mu}$ is the photon frequency measured in the plasma frame. $B^{\mu}_{\bot}$ represents the spatial magnetic field in the plasma frame with the form
$B^{\mu}_{\bot}=B^{\mu}-\omega^{-2}(B\cdot p)p^{\mu}_{\bot},$ (5)
where $p^{\mu}_{\bot}$ is the spatial four-momentum of photon,
$p^{\mu}_{\bot}=p^{\mu}-\omega U^{\mu}.$ (6)
The celestial coordinates $(\alpha,\beta)$ of a photon emitted by plasma synchrotron radiation in the screen of a distant observer are related to its conserved quantities $(\lambda,\eta)$ by Himwich:2020msm ; Bardeen:1973tla
$\alpha=-\frac{\lambda}{\sin\theta_{0}},\quad\quad\quad\beta=\pm_{o}\sqrt{\eta+a^{2}\cos^{2}\theta_{0}-\lambda^{2}\cot^{2}\theta_{o}},$ (7)
where $\theta_{o}$ represents the observational angle and $\pm_{o}$ denotes the sign of $p_{\theta}$ at the observer. With the components of $f^{\mu}$ projected along the $\alpha$ and $\beta$ directions on the observer’s screen, one can obtain the electric vector polarization angle (EVPA)
$\chi=\tan^{-1}{\frac{f^{\beta}}{f^{\alpha}}}\,.$ (8)
Since the geometry of a rotating hairy Horndeski black hole (1) belongs to a family of Petrov Type $D$ spacetimes, the Penrose-Walker quantity $\kappa$ Walker:1970un ; Chandrasekhar:1985kt
$\kappa=\kappa_{1}+i\kappa_{2}=2\Psi_{2}^{-\frac{1}{3}}p^{\mu}f^{\nu}\left(l_{[\mu}n_{\nu]}-m_{[\mu}\bar{m}_{\nu]}\right)\,$, (9)
is conversed along null geodesics. $l_{\mu}$, $n_{\mu}$, $m_{\mu}$ and $\bar{m}_{\mu}$ are vectors in the null basis $(l,n,m,\bar{m})$. With the four-moment of photon $p^{\mu}$ and the polarization vector $f^{\mu}$, the Penrose-Walker constant in the rotating hairy Horndeski black hole spacetime (1) can be expressed as
$\kappa=\kappa_{1}+i\kappa_{2}=(A-iB)\Psi_{2}^{-\frac{1}{3}},$ (10)
$A=2(p^{[t}f^{r]}+a\sin^{2}\theta p^{[t}f^{\phi]}),$ (11)
$B=2\sin{\theta}\left[(r^{2}+a^{2})p^{[\phi}f^{\theta]}-ap^{[t}f^{\phi]}\right]\,$, (12)
where $\kappa_{1}$ and $\kappa_{2}$ represent the real and imagine part of $\kappa$. The non-vanishing Weyl scalar $\Psi_{2}$ has a form
$\Psi_{2}=-\frac{1}{2(r-ia\cos\theta)^{2}}\bigg{[}\frac{2M-h}{(r-ia\cos\theta)}+\frac{h(5r+ia\cos\theta)}{6r(r+ia\cos\theta)}\bigg{]}.$ (13)
As in ref.Hou:2024qqo , to analyze the near-horizon features of the polarizations of emissions from the plasma on the equatorial plane, one can assume that the flow’s velocity component $U^{\theta}=0$, while the covariant components $U_{t}$ and $U_{\phi}$ keep invariant near the black hole horizon. The main reason is that the accretion lies on the equatorial plane and the near-horizon plasma flow is dominated by gravitational attraction, which means that the $U_{t}$ and $U_{\phi}$ are finite and change slowly in the near-horizon region. This assumption is supported by the simulated results of MADs presented in Chael:2021rjo ; Ricarte:2022wpd . With the above assumption, we find that the four-velocity of the equatorial accreted flow near the outer horizon of a rotating hairy Horndeski black hole (1) can be expressed as
$U^{t}$ $=$ $-\frac{4(1-\frac{h}{2}\ln{\frac{r_{+}}{2}})^{2}(U_{t}+\Omega_{h}U_{\phi})}{\Delta}-\bigg{[}1+\frac{2}{r_{+}}\bigg{(}1+h-\frac{h}{2}\ln\frac{r_{+}}{2}-\frac{h(2r_{+}+h)}{2(r_{+}-1+\frac{h}{2}+\frac{h}{2}\ln\frac{r_{+}}{2})}\bigg{)}\bigg{]}U_{t}$ (14)
$+$ $\frac{aU_{\phi}(1+\frac{h}{2}-\frac{h}{2}\ln{\frac{r_{+}}{2}})}{r_{+}^{2}(r_{+}-1+\frac{h}{2}+\frac{h}{2}\ln\frac{r_{+}}{2})}+o(\Delta^{0}),$
$U^{r}$ $=$ $\frac{-2|(1-\frac{h}{2}\ln{\frac{r_{+}}{2}})(U_{t}+\Omega_{h}U_{\phi})|}{r_{+}}+o(\Delta^{0}),$ (15)
$U^{\theta}$ $=$ $0,$ (16)
$U^{\phi}$ $=$ $-\frac{2a(1-\frac{h}{2}\ln{\frac{r_{+}}{2}})(U_{t}+\Omega_{h}U_{\phi})}{\Delta r_{+}}+\frac{(1+\frac{h}{2}-\frac{h}{2}\ln\frac{r_{+}}{2})(U_{\phi}+aU_{t})}{r_{+}^{2}(r_{+}-1+\frac{h}{2}+\frac{h}{2}\ln\frac{r_{+}}{2})}+o(\Delta^{0})\,$, (17)
where $\Omega_{h}=\frac{a}{r^{2}_{+}+a^{2}}$ is the angular velocity at the outer horizon. Obviously, as $h\rightarrow 0$, the four-velocity of the equatorial accreted flow $U^{\mu}$ tends to that in the Kerr case obtained in ref.Hou:2024qqo . At the outer horizon, from $\Delta(r_{+})=0$, one has $r^{2}_{+}+a^{2}=2r_{+}(1-\frac{h}{2}\ln\frac{r_{+}}{2})>0$, which means that $U^{r}<0$ for the accreting plasma as in the Kerr black hole case. Inserting Eqs.(14)-(17) into Eq. (3), one can obtain the near-horizon expansion of the magnetic field
$B^{t}=\frac{2\Psi U_{t}(1-\frac{h}{2}\ln\frac{r_{+}}{2})}{r_{+}\Delta}\frac{(U_{t}+\Omega_{h}U_{\phi})}{|U_{t}+\Omega_{h}U_{\phi}|}+o(\Delta^{-1}),$ (18)
$B^{r}=\frac{\Psi}{r_{+}^{2}}U_{t},\quad\quad\quad B^{\theta}=0,$ (19)
$B^{\phi}=\frac{\Psi U_{t}a}{r_{+}^{2}\Delta}\frac{(U_{t}+\Omega_{h}U_{\phi})}{|U_{t}+\Omega_{h}U_{\phi}|}+o(\Delta^{-1})\,.$ (20)
Similarly, one can obtain $B^{\theta}=0$ from $U^{\theta}=0$. Moreover, for the plunging matter, one has $U_{t}+\Omega_{h}U_{\phi}<0$ since a probe particle with $U_{t}+\Omega_{h}U_{\phi}>0$ cannot classically reach the event horizon. For the roating hairy Horndeski black hole, we find that the ratio also meets $B^{\phi}/B^{r}\approx-a/\Delta$, which agrees with that from the simulation presented in Ricarte:2022wpd . Expanding the polarization vector $f^{\mu}$ near the outer horizon $r=r_{+}$, one has
$\frac{X}{\varepsilon}f^{t}=\frac{-a(a^{2}+r_{+}^{2})\sigma_{\theta}\sqrt{\eta}}{r_{+}^{2}\Delta}+\frac{a\sigma_{\theta}\sqrt{\eta}}{r_{+}^{2}}+\frac{2a^{3}\sigma_{\theta}\sqrt{\eta}}{r_{+}^{3}(h+2r_{+}+h\ln{\frac{r_{+}}{2}})}+o(\Delta^{0})$ (21)
$\frac{X}{\varepsilon}f^{r}=\frac{a\sigma_{\theta}\sqrt{\eta}}{r_{+}^{2}}+R_{1}\Delta+o(\Delta)$ (22)
$\frac{X}{\varepsilon}f^{\theta}=\frac{a(a\lambda-a^{2}-r_{+}^{2})-\sigma_{r}a|a\lambda-a^{2}-r_{+}^{2}|}{r_{+}^{2}\Delta}+\Theta_{1}+o(\Delta^{0})$ (23)
$\frac{X}{\varepsilon}f^{\phi}=-\frac{a^{2}\sigma_{\theta}\sqrt{\eta}}{r_{+}^{2}\Delta}+\frac{\sigma_{\theta}\sqrt{\eta}}{r_{+}^{2}}\frac{2a^{3}\sigma_{\theta}\sqrt{\eta}}{r_{+}^{3}(h+2r_{+}+h\ln{\frac{r_{+}}{2}})}+o(\Delta^{0})\,$, (24)
where $X=-\Psi^{-1}\sqrt{-g(U^{\mu}p_{\mu})B^{2}_{\bot}}$, $\varepsilon=-p_{t}$, $\lambda$ and $\eta$ are conserved impact parameters for the emitted photons. The term $\sigma_{r}$ and $\sigma_{\theta}$ denotes the sign of $p^{r}$ and $p^{\theta}$, respectively. The forms of $R_{1}$ and $\Theta_{1}$ are rather complicated and we do not list them here.
Substituting Eqs.(21)-(24) into Eq.(9), we can obtain the expansion of the Penrose-Walker constant $\kappa$ near outer horizon of the rotating hairy Horndeski black hole, and find that the ratio
$z=\frac{\kappa_{1}}{\kappa_{2}}=z_{0}+z_{1}\Delta+o(\Delta^{2}),$ (25)
is independent of the plasma motion near the horizon as in the Kerr case Hou:2024qqo , where the coefficients $z_{0}$ and $z_{1}$ can be expressed respectively as
$z_{0}=\frac{\sigma_{\theta}\sqrt{\eta}}{\lambda-a},\quad\quad\quad z_{1}=\frac{\sigma_{\theta}\sqrt{\eta}(1+z^{2}_{0})\Omega_{h}}{2a^{2}(\Omega_{h}\lambda-1)}.$ (26)
Obviously, the forms of the coefficients $z_{0}$ and $z_{1}$ coincide with those in the Kerr case. However, we find that in the rotating hairy Horndeski black hole case, for fixed $\lambda$ and $\eta$, the coefficient $z_{0}$ depends on only the black hole spin parameter $a$, while the coefficient $z_{1}$ is related to the hairy parameter $h$ by the angular velocity $\Omega_{h}$ of the outer horizon. According to Ref.Himwich:2020msm , the EVPA on the observer’s plane can be derived from Eq.(8)and Eq.(25) ,and is given by
$\chi=\tan^{-1}\left(\frac{\beta z+\mu}{\beta-\mu z}\right)=\tan^{-1}\left(\frac{\mu}{\beta}\right)+\tan^{-1}z=\tan^{-1}\left(\frac{\mu}{\beta}\right)+\tan^{-1}z_{0}+\frac{z_{1}\Delta}{1+z_{0}^{2}}+o(\Delta^{1}),$ (27)
where $\mu=-(\alpha+a\sin{\theta_{o}})$. From Eqs.(25), (26) and (27), it is found that the hairy parameter $h$ appears explicitly only in the sub-leading term. However, the effects of $h$ on the observed EVPA $\chi$ also include contributions from the leading terms because the parameters $\lambda$ and $\eta$ of photons traveling from region near-horizon to a specific observer are not arbitrary, instead, they should be certain specific values related to the hairy parameter $h$. This means that the polarization structures near horizon could provide a potential tool to detect black hole’s hairs.
Figure 3: Variations of $z$ with $\Delta$ for three different types of plasma flows in the rotating hairy Horndeski black hole spacetime. The left and right panels correspond to spin parameters $a=0.2$ and $a=0.94$, respectively. Each panel in the bottom row is a partial enlargement of the respective figure in the upper row.
III Impacts of the Hairy Parameter on the Near-Horizon Polarized Image of a rotating Horndeski Black Hole
Let us now to discuss features of the near-horizon polarized image of a rotating Horndeski black hole (1) and to probe the corresponding impacts of the hairy parameter $h$. In Fig.(3), we present variations of the ratio $z$ with the quantity $\Delta$ in the rotating hairy Horndeski black hole spacetime for three different plasma flow conditions as in Hou:2023bep . The plasma is assumed to infall from infinity in the type I model and to plunge from the prograde innermost stable circular orbit (ISCO) in the type II model, respectively. The type III model corresponds to the fitting result from the simulation of MAD introduced in Chael:2021rjo . Our results show that the curves $z-\Delta$ are almost overlapped in the range $\Delta<10^{-2}$ for all three types of plasma flows in the black hole spacetimewith different values of the hairy parameter $h$. This further confirms that effects arising from types of plasma flow on the polarization structures near black hole horizon are negligible. Furthermore, from their partial enlargement, we find that the curves $z-\Delta$ are separate for different hairy parameter $h$, while still are overlapped for above three types of plasma flows. This means that the effects from the hairy parameter $h$ is much stronger than those from the types of plasma flows. In additional, the ratio $z$ increases with the hairy parameter $h$.
Figure 4: Variation of the EVPAs in the near horizon images with the azimuthal angle $\phi$ in the rotating hairy Horndeski black hole spacetime. The left panel is for $\theta_{o}=17^{\circ}$ and the right one is for $\theta=80^{\circ}$. In each panel, the solid and dashed lines correspond to the cases $h=h_{c}$ and $h=0$, respectively. The blue and red lines denote the cases $a=0.94$ and $a=0.2$, respectively.
In Fig.4, we present the changes of the observed EVPA of polarization vectors with the the azimuth angle $\phi$ for the photons from region near the event horizon for a rotating hairy Horndeski black hole with different hairy parameter $h$. As expected, the observed EVPA of polarization vectors significantly changes with the black hole’s spin $a$ and the observed inclination $\theta_{o}$. Furthermore, we find that the effects of hairy parameter $h$ on the observed EVPA $\chi$ depend on the black hole’s spin $a$ and the observed inclination $\theta_{o}$. For a fixed inclination $\theta_{o}$, in the slowly rotating case where $a=0.2$, the observed EVPA $\chi$ corresponding to $h=h_{c}$ is greater than that associated with $h=0$. The situation is reversed in the rapidly rotating case with $a=0.94$. For a fixed black hole spin, as the azimuthal angle $\phi$ is in the range $(3,6)$, one can find that the impact of $h$ on the observed EVPA is more distinct in the low inclination case with $\theta_{o}=17^{\circ}$. As the azimuthal angle $\phi$ is beyond the range $(3,6)$, the difference between the observed EVPAs induced by the hairy parameter $h$ is almost independent of the observer inclination $\theta_{o}$. Therefore, the observer inclination’s impact on the EVPA difference caused by the hairy parameter $h$ is also shaped by the azimuthal angle of the observed polarization vector. Through making a comparison, We also find that the dependence of the hairy parameter’s impact on the EVPA is stronger with respect to the black hole spin $a$ than to the inclination angle $\theta_{o}$.
Figure 5: The EVPAs in the near horizon images of rotating hairy Horndeski black holes are represented by different symbols: $\bullet$ indicates $\chi=\frac{\pi}{6}$, and $\triangle$ indicates $\chi=-\frac{\pi}{6}$. The left panel is for $a=0.2$ and the right one is for $a=0.94$. In each panel, the curves represent the near horizon images. The solid and dashed lines correspond to the cases $h=h_{c}$ and $h=0$, respectively. The blue and red lines denote the cases $\theta_{o}=17^{\circ}$ and $\theta_{o}=80^{\circ}$, respectively.
In Fig.5, we present the positions of the polarization vectors related to EVPA $\chi=\frac{\pi}{6}$ or $\chi=-\frac{\pi}{6}$ in the near horizon images for a rotating hairy Horndeski black hole with spin parameters $a=0.2$ and $a=0.94$, observed at inclination angles $\theta_{o}=17^{\circ}$ and $\theta_{o}=80^{\circ}$. It is show that the position of polarization vectors with the same EVPA $\chi=\frac{\pi}{6}$ or $\chi=-\frac{\pi}{6}$ changes with the hairy parameter $h$. Apart from the size of the event horizon image, the variation in the position of polarization vectors with the same EVPA, induced by the hairy parameter $h$, is most strongly affected by the black hole spin. Although the effects from the hairy parameter $h$ in the near horizon images is weaker than those from black hole spin and observed inclination, one can infer the information on black hole’s hair by integrating the corresponding observations with higher precision.
IV Discussions and conclusions
We have investigated the observed polarization in the near-horizon images of a rotating hair black hole within Horndeski gravity. Besides its mass and spin parameter, this black hole possesses an additional hairy parameter $h$, which originates from Horndeski theory. Our results confirm once again that the EVPA of polarization vectors in the near-horizon images is governed exclusively by spacetime geometry and is irrespective of the plasma flow geometry as found in Hou:2024qqo . Furthermore, the effects of the hairy parameter $h$ on the observed EVPA depend on the black hole’s spin $a$ and the observed inclination $\theta_{o}$. For a fixed inclination $\theta_{o}$, in the slowly rotating case where $a=0.2$, the observed EVPA corresponding to $h=h_{c}$ is greater than that associated with $h=0$. The situation is reversed in the rapidly rotating case with $a=0.94$. For a fixed black hole spin, the dependence of the difference between the observed EVPAs induced by the hairy parameter $h$ on the observer inclination $\theta_{o}$ is also influenced by the azimuthal angle of the observed polarization vector. As the azimuthal angle $\phi$ is in the range $(3,6)$, the impact of $h$ on the observed EVPA is more distinct in the low inclination case with $\theta_{o}=17^{\circ}$. However, as the azimuthal angle $\phi$ is beyond the range $(3,6)$, the difference between the observed EVPAs induced by the hairy parameter $h$ is almost independent of the observer inclination $\theta_{o}$. Furthermore, the dependence of the hairy parameter’s impact on the EVPA is stronger with respect to the black hole spin $a$ than to the inclination angle $\theta_{o}$. Although the effects from the hairy parameter $h$ in the near horizon images is weaker than those from black hole spin and the observed inclination, one can infer the information on black hole’s hair by integrating the corresponding observations with higher precision, which could help to further understand the no-hair theorem.
Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2020YFC2201400) and National Natural Science Foundation of China (Grant Nos. 12275079 and 12035005), and the innovative research group of Hunan Province under Grant No. 2024JJ1006.